Project

General

Profile

Actions

Location : CR21 at 14:00 pm

Phone access : New phonebridge number
UK: 08082380274
US: 631.621.5253
Access code: 1779339517

Agenda

  • Actions
    • Macwaters - Remove old overheated CSW71 to R9 waiting for John Govans and the van - in progress
    • Hall Manager - water system change-over procedure (stands) - have prescription, need document (Colin?)
    • Progress on Hall list attachment:MICE-Work-List-May-2017.pdf (Colin) Remaining items_
    • CWhyte to request survey of CC Magnets after cycle - stands
    • Tracker group: decide what should be done to prevent repeat contamination of Tracker Cryostats
    • Status of miceraid5 disk cleanup (stands) EOverton now has a script but needs to run it: needs to hand over the procedure and script to someone, i.e. Durga
  • DC update
  • Hydrogen system update and plans
  • Issues to fix before September : Cycle 2017 Issues : This will be discussed after a prioritisation exercise is carried out.
  • Mobile Logistics and Operational Support update (CWhyte) (If I'd known this was about steak frites I'd have put it further up the agenda! SB)
  • AOB

Minutes

  • Actions
    • Macwaters - Remove old overheated CSW71 to R9 waiting for John Govans and the van - CSW71 in Tracker lab, F70 returned to lender - done.
    • Hall Manager - water system change-over procedure (stands) - have prescription and draft document (Colin/Josef)
    • Progress on Hall list attachment:MICE-Work-List-May-2017.pdf (Colin) Last remaining item is believed to be verifying the interlocks; hope to get Trevor Hartnett down in next few weeks?
    • CWhyte to request survey of CC Magnets after cycle: need survey of SS's (Tracker positions during data taking) and FC (Absorber location) before pulling the FC out - stands
    • Tracker group: decide what should be done to prevent repeat contamination of Tracker Cryostats
      Need to decide if warming up and drying cryostats again over the summer. Decision to be made in tomorrow's Tracker meeting.
    • Status of miceraid5 disk cleanup (stands) EOverton now has a script but needs to run it: needs to hand over the procedure and script to someone, i.e. Durga stands
  • DC update
    FC Quench test last Wed.
    DS fridge has been serviced by Joerg. Still have 8 bar in buffer tank.
    PRY is open, waiting for FC to warm up before removing it. MTucker to let Nitrogen into some of the vacuum spaces later today.
    Another widespread Watchdog Timer failure in the CANbuses last Thursday. This time it also affected the Target interlocks and the DS vacuum; in addition to the cryo-compressors and general vacuum like last time. Is Weird Shit since these CANbuses are not interlinked. Possibly co-incides with MT cutting power off to A/C 5 - DL have been asked to check the Power Quality Monitor (but Target Vac-Rack on a different substation?). Need to consider what states things go to when this error occurs.
    AKurup has restarted the State Machines and now needs to re-set the Alarm Handlers to get everything consistent.
  • SS-PRY Spacers
    JTarrant is checking that we have the correct load-case. Need to finish the modelling, then decide if we need them by the September run. CWhyte to close out with JT.
  • Hydrogen system update and plans
    Plan: Tue/Wed. final changes to turret pipework in R9, then need to check over and get the radiation shield and thermal links fitted. Then pull FC out and extract turret (subject to survey timing).
    Have shown that we can control liquid Neon stably. Have identified some needed control software changes; Phil implementing at DL then needs installing at RAL (PW to RAL).
    Need to repeat the Neon test with the new turret and software (order more Neon).
    LH2 Controls: plan is to have the controls PLC send data to a gateway machine which distributes the PVs on micenet, will need sign-off from Bob Mannix. Also need to check with Chris Brew (Village Manager) that the implementation won't mess up other things on micenet.
  • Mobile Logistics and Operational Support update (CWhyte)
    No.
  • AOB
    PF has tried to swap the IP numbers of A/C units 2 and 3, but we don't have the web UI password
    It's not yet known when John Govans will be back.

Updated by Nebrensky, Henry almost 6 years ago · 16 revisions